On the twin designs with the Ionin–type parameters

H. Kharaghani^{∗} Department of Mathematics & Computer Science University of Lethbridge Lethbridge, Alberta, T1K 3M4 Canada hadi@cs.uleth.ca

Submitted: July 31, 1999; Accepted: October 20, 1999

Dedicated to Professor Reza Khosrovshahi on the occasion of his 60th birthday

Keywords: Symmetric design, regular Hadamard matrix, Bush-type Hadamard matrix, design with Ionin–type parameters, balanced generalized weighing matrix, weighing matrix. MR Subject Code: 05B05

Abstract

Let 4n^{2} be the order of a Bush-type Hadamard matrix with q = (2n − 1)^{2} a prime power. It is shown that there is a weighing matrix

W (4(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, 4q^{m}n^{2})

which includes two symmetric designs with the Ionin–type parameters

ν = 4(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, κ = q^{m}(2n^{2} − n), λ = q^{m}(n^{2} − n)

for every positive integer m. Noting that Bush–type Hadamard matrices of order 16n^{2} exist for all n for which an Hadamard matrix of order 4n exist, this provides a new class of symmetric designs.

∗Thanks to Stephen Ney for proving me wrong on my first choice of the cyclic group by writing a program and applying the group. Wolfgang Holzmann, as always, was a great helper.

1 Introduction

Recently Ionin [3] introduced an elegant method to use a very special class of regular Hadamard matrix of order 36 in a class of balanced generalized weighing matrices to construct a large class of symmetric designs. The key to his construction is the existence of a class of balanced generalized weighing matrices BGW (q^{m}+q^{m−1}+⋯+q+1, q^{m}, q^{m}− q^{m−1}) over a cyclic group of order t, where q is a prime power, m is a positive integer and t is a divisor of q − 1. A balanced generalized weighing matrix BGW (ν, κ, λ) over a group G is a matrix W = [w_{ij}] of order ν , with w_{ij} ∈ G ∪ {0} such that each row and column of W has κ non–zero entries and for each k ≠ l, the multiset {w_{kj}w_{lj}^{−1} : 1 ≤ j ≤ ν, w_{kj} ≠ 0, w_{lj} ≠ 0} contains λ/|G| copies of every element of G. For his construction, Ionin starts with what he calls a starting regular Hadamard matrix. He then constructs a cyclic group from certain one to one maps on the starting Hadamard matrix. The final step is a clever use of the above known balanced generalized weighing matrices and the starting regular Hadamard matrix. See [3] for details.

In this paper we use a Bush-type Hadamard matrix. A Bush-type Hadamard matrix is a block matrix H = [H_{ij}] of order 4n^{2} with block size 2n, H_{ii} = J_{2n} and H_{ij}J_{2n} = J_{2n}H_{ij} = 0, i ≠ j, 1 ≤ i ≤ 2n, 1 ≤ j ≤ 2n, where J_{2n} is the 2n by 2n matrix of all entries 1. We then introduce a cyclic group consisting of signed permutation matrices of order 4n^{2}. Using the above balanced generalized weighing matrices over this cyclic group and a Bush-type Hadamard matrix, we will construct a very special weighing matrix (a weighing matrix is a (0, ±1)-matrix with orthogonal rows and columns). By replacing certain blocks of this weighing matrix with zero blocks we get a (0, ±1)-matrix which becomes a symmetric design with Ionin–type parameters;

ν = 4(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, κ = q^{m}(2n^{2} − n), λ = q^{m}(n^{2} − n),

by replacing either all 1’s or all −1’s with zeros. The construction method in this paper works for any Bush–type Hadamard matrix and is potentially very useful for constructing many other designs. One of the differences between the construction here and that of Ionin is the way that the cyclic group being used. The cyclic group used in this paper acts as a multiplication by signed permutation matrix, while Ionin’s cyclic group consists of column permutations. This makes our computations easier than that of Ionin.

Since Bush-type Hadamard matrices of order 16n^{2} is known to exist if 4n is the order of an Hadamard matrix, we get many more designs than Ionin. However, Ionin’s method [3] is essential for orders 4n^{2} when n is an odd integer. The problem of investigating the existence of Bush–type Hadamard matrices of order 4n^{2}, n an odd integer, is a tough one. It is quite interesting to note that if such matrices exist, then the construction method given in this paper simplifies Ionin’s method substantially. On the other hand the non–existence of these matrices would imply the non–existence of projective planes of order 2n, n an odd integer. The smallest order for which the existence of Bush–type Hadamard matrices is unknown is 36. We conjecture here that Bush-type Hadamard matrices exist for all orders 4n^{2}, n an odd integer.

For a (0, ±1)-matrix K, let K = K^{+} − K^{−}, where K^{+} and K^{−} are (0, 1)-matrices. The Kronecker product of two matrices A = [a_{ij}] and B, denoted A ⊗ B is defined, as usual, by A ⊗ B = [a_{ij}B]. Throughout the paper P is the block matrix defined by P = (J_{2n}−I_{2n})⊗J_{2n}, where I_{2n} is the identity matrix of order 2n. For a matrix A = [a_{ij}], denote by |A| the matrix [|a_{ij}|]. Throughout the paper − represent −1.

2 Bush-type Hadamard matrices and twin designs

K. A. Bush [1] proved that if there exists a projective plane of order 2n, then there is an Hadamard matrix H of order 4n^{2}, such that:

1. H is symmetric.

2. H = [H_{ij}], where H_{ij} are blocks of order 2n, H_{ii} = J_{2n} and H_{ij}J_{2n} = J_{2n}H_{ij} = 0, for i ≠ j, 1 ≤ i ≤ 2n, 1 ≤ j ≤ 2n.

Bush’s interest was mainly in the non-existence of such matrices. While there are different methods to construct matrices of above type of order 16n^{2}, (see [5]), we are not aware of non–existence of matrices of this form of order 4n^{2} for a single odd value of n, n > 1.

Please note that in this paper by a Bush-type Hadamard matrix we mean an Hadamard matrix satisfying only condition 2 above, as we do not need to assume that H is symmetric for our construction. For completeness we include the following result of the author [5].

Theorem 1 Let 4n be the order of an Hadamard matrix, then there is a Bush-type Hadamard matrix of order 16n^{2}.

Proof.

Let K be a normalized Hadamard matrix of order 4n. Let r_{1}, r_{2}, …, r_{4n} be the row vectors of K. Let C_{i} = r_{i}^{t}r_{i}, i = 1, 2, …, 4n. Then it is easy to see that:

1. C_{i}^{t} = C_{i}, for i = 1, 2, …, 4n.

2. C_{1} = J_{4n}, C_{i}J_{4n} = J_{4n}C_{i} = 0, for i = 2, …, 4n. the electronic journal of combinatorics 7 (2000), #R1 4

3. C_{i}C_{j}^{t} = 0, for i ≠ j, 1 ≤ i, j ≤ 4n.

4. ∑_{i=1}^{4n} C_{i}C_{i}^{t} = 16n^{2}I_{4n}. Now let H = circ(C_{1}, C_{2}, …, C_{4n}), the block circulant matrix with first row C_{1}C_{2} … C_{4n}. Then H is a Bush-type Hadamard matrix of order 16n^{2}. Example 2 Let

K = ( \begin{matrix}1 & 1 & 1 & 1 \\ 1 & 1 & − & − \\ 1 & − & 1 & − \\ 1 & − & − & 1 \end{matrix} )

Then,

r_{1} = ( 1 1 1 1 )

r_{2} = ( 1 1 − − ) r_{3} = ( 1 − 1 − ) r_{4} = ( 1 − − 1 )

C_{1} = r_{1}^{t}r_{1} = ( \begin{matrix}1 & 1 & 1 & 1 \\ 1 & 1 & 1 & 1 \\ 1 & 1 & 1 & 1 \\ 1 & 1 & 1 & 1 \end{matrix} )

C_{2} = r_{2}^{t}r_{2} = ( \begin{matrix}1 & 1 & − & − \\ 1 & 1 & − & − \\ − & − & 1 & 1 \\ − & − & 1 & 1 \end{matrix} )

C_{3} = r_{3}^{t}r_{3} = ( \begin{matrix}1 & − & 1 & − \\ − & 1 & − & 1 \\ 1 & − & 1 & − \\ − & 1 & − & 1 \end{matrix} )

C_{4} = r_{4}^{t}r_{4} = ( \begin{matrix}1 & − & − & 1 \\ − & 1 & 1 & − \\ − & 1 & 1 & − \\ 1 & − & − & 1 \end{matrix} )

Then H = circ(C_{1}, C_{2}, C_{3}, C_{4}) is the following matrix,

[ \begin{matrix}1 & 1 & 1 & 1 & 1 & 1 & − & − & 1 & − & 1 & − & 1 & − & − & 1 \\ 1 & 1 & 1 & 1 & 1 & 1 & − & − & − & 1 & − & 1 & − & 1 & 1 & − \\ 1 & 1 & 1 & 1 & − & − & 1 & 1 & 1 & − & 1 & − & − & 1 & 1 & − \\ 1 & 1 & 1 & 1 & − & − & 1 & 1 & − & 1 & − & 1 & 1 & − & − & 1 \\ 1 & − & − & 1 & 1 & 1 & 1 & 1 & 1 & 1 & − & − & 1 & − & 1 & − \\ − & 1 & 1 & − & 1 & 1 & 1 & 1 & 1 & 1 & − & − & − & 1 & − & 1 \\ − & 1 & 1 & − & 1 & 1 & 1 & 1 & − & − & 1 & 1 & 1 & − & 1 & − \\ 1 & − & − & 1 & 1 & 1 & 1 & 1 & − & − & 1 & 1 & − & 1 & − & 1 \\ 1 & − & 1 & − & 1 & − & − & 1 & 1 & 1 & 1 & 1 & 1 & 1 & − & − \\ − & 1 & − & 1 & − & 1 & 1 & − & 1 & 1 & 1 & 1 & 1 & 1 & − & − \\ 1 & − & 1 & − & − & 1 & 1 & − & 1 & 1 & 1 & 1 & − & − & 1 & 1 \\ − & 1 & − & 1 & 1 & − & − & 1 & 1 & 1 & 1 & 1 & − & − & 1 & 1 \\ 1 & 1 & − & − & 1 & − & 1 & − & 1 & − & − & 1 & 1 & 1 & 1 & 1 \\ 1 & 1 & − & − & − & 1 & − & 1 & − & 1 & 1 & − & 1 & 1 & 1 & 1 \\ − & − & 1 & 1 & 1 & − & 1 & − & − & 1 & 1 & − & 1 & 1 & 1 & 1 \\ − & − & 1 & 1 & − & 1 & − & 1 & 1 & − & − & 1 & 1 & 1 & 1 & 1 \end{matrix} ]

Lemma 3 Let H = [H_{ij}] be a Bush-type Hadamard matrix of order 4n^{2}. Let M = H − I_{2n} ⊗ J_{2n}. Then each of M^{+} and M^{−} is a symmetric (4n^{2}, 2n^{2} − n, n^{2} − n)–design.

Proof.

The row sums of H are all 2n. Thus the negative entries in H (hence in M) can be viewed as the incidence matrix of a symmetric (4n^{2}, 2n^{2} − n, n^{2} − n)-design. Since negating all the off diagonal blocks of H leaves it Bush-type Hadamard, the positive entries of M form another symmetric (4n^{2}, 2n^{2} − n, n^{2} − n)-design.

Alternate Proof.

We need this proof for use in the proof of our main result. First note that,

M = M^{+} − M^{−},
P = M^{+} + M^{−}.

So, 2M^{+} = M + P. Now it is easy to see that, MP^{t} = P M^{t} = 0 and thus 4M^{+}M^{+t} = MM^{t} + P P^{t} and M^{+}M^{+t} = M^{−}M^{−t}. Now,

MM^{t} + P P^{t} = (H − I_{2n} ⊗ J_{2n})(H^{t} − I_{2n} ⊗ J_{2n}) + ((J_{2n} − I_{2n}) ⊗ J_{2n})((J_{2n} − I_{2n}) ⊗ J_{2n})
= HH^{t} − 2nI_{2n} ⊗ J_{2n} + 4n(n − 1)J_{2n} ⊗ J_{2n} + 2nI_{2n} ⊗ J_{2n}
= 4n^{2}I_{4n^{2}} + 4n(n − 1)J_{2n} ⊗ J_{2n}.

Therefore, M^{+}M^{+t} = M^{−}M^{−t} = n^{2}I_{4n^{2}} + n(n − 1)J_{2n} ⊗ J_{2n}. This means that both M^{+} and M^{−} are symmetric (4n^{2}, 2n^{2} − n, n^{2} − n)–designs.

Note that the two designs may not be isomorphic in general (and the problem of finding out when the two are equivalent is not an easy one). Following [4] We call the matrix M a twin design.

Example 4

Let C_{2}, C_{3}, C_{4} be the matrices of the previous example. Then M = circ(0, C_{2}, C_{3}, C_{4}) is the following matrix,

[ \begin{matrix}0 & 0 & 0 & 0 & 1 & 1 & − & − & 1 & − & 1 & − & 1 & − & − & 1 \\ 0 & 0 & 0 & 0 & 1 & 1 & − & − & − & 1 & − & 1 & − & 1 & 1 & − \\ 0 & 0 & 0 & 0 & − & − & 1 & 1 & 1 & − & 1 & − & − & 1 & 1 & − \\ 0 & 0 & 0 & 0 & − & − & 1 & 1 & − & 1 & − & 1 & 1 & − & − & 1 \\ 1 & − & − & 1 & 0 & 0 & 0 & 0 & 1 & 1 & − & − & 1 & − & 1 & − \\ − & 1 & 1 & − & 0 & 0 & 0 & 0 & 1 & 1 & − & − & − & 1 & − & 1 \\ − & 1 & 1 & − & 0 & 0 & 0 & 0 & − & − & 1 & 1 & 1 & − & 1 & − \\ 1 & − & − & 1 & 0 & 0 & 0 & 0 & − & − & 1 & 1 & − & 1 & − & 1 \\ 1 & − & 1 & − & 1 & − & − & 1 & 0 & 0 & 0 & 0 & 1 & 1 & − & − \\ − & 1 & − & 1 & − & 1 & 1 & − & 0 & 0 & 0 & 0 & 1 & 1 & − & − \\ 1 & − & 1 & − & − & 1 & 1 & − & 0 & 0 & 0 & 0 & − & − & 1 & 1 \\ − & 1 & − & 1 & 1 & − & − & 1 & 0 & 0 & 0 & 0 & − & − & 1 & 1 \\ 1 & 1 & − & − & 1 & − & 1 & − & 1 & − & − & 1 & 0 & 0 & 0 & 0 \\ 1 & 1 & − & − & − & 1 & − & 1 & − & 1 & 1 & − & 0 & 0 & 0 & 0 \\ − & − & 1 & 1 & 1 & − & 1 & − & − & 1 & 1 & − & 0 & 0 & 0 & 0 \\ − & − & 1 & 1 & − & 1 & − & 1 & 1 & − & − & 1 & 0 & 0 & 0 & 0 \end{matrix} ]

3 A cyclic subgroup of signed permutation matrices

Let SP_{m} be the set of all signed permutation matrices of order m. Let U = circ(0, 1, 0, …, 0) be the circulant shift permutation matrix of order 2n (this is a circulant matrix of order 2n with first row 010 … 0) and N = diag(−, 1, 1, …, 1) be the diagonal matrix of order 2n with −1 at the (1, 1)-position and 1 elsewhere on the diagonal. Let E = UN, then E is in SP_{2n}. Let G_{4n} = {γ^{i} = E^{i} ⊗ I_{2n}: i = 1, 2, …, 4n} =≺ γ ≻.

Lemma 5 G_{4n} is a cyclic subgroup of SP_{4n^{2}} of order 4n.

Proof.

For 1 ≤ r ≤ 2n, (UN)^{r} is U^{r} with its first r columns negated. Thus γ^{2n} = E^{2n} ⊗ I_{2n} = −I_{2n} ⊗ I_{2n} = −I_{4n^{2}}. It follows now that G_{4n} is a cyclic subgroup of SP_{4n^{2}} of order 4n.

Note that G_{4n} is a (signed) group subgroup of SP_{4n^{2}} and ∑_{γ∈G_{4n}} γ = 0.

Example 6

Let n = 2 in lemma 5,

γ = E ⊗ I_{4} = ( \begin{matrix}0 & I_{4} & 0 & 0 \\ 0 & 0 & I_{4} & 0 \\ 0 & 0 & 0 & I_{4} \\ −I_{4} & 0 & 0 & 0 \end{matrix} )

γ^{2} = E^{2} ⊗ I_{4} = ( \begin{matrix}0 & 0 & I_{4} & 0 \\ 0 & 0 & 0 & I_{4} \\ −I_{4} & 0 & 0 & 0 \\ 0 & −I_{4} & 0 & 0 \end{matrix} )

γ^{3} = E^{3} ⊗ I_{4} = ( \begin{matrix}0 & 0 & 0 & I_{4} \\ −I_{4} & 0 & 0 & 0 \\ 0 & −I_{4} & 0 & 0 \\ 0 & 0 & −I_{4} & 0 \end{matrix} )

γ^{4} = E^{4} ⊗ I_{4} = ( \begin{matrix}−I_{4} & 0 & 0 & 0 \\ 0 & −I_{4} & 0 & 0 \\ 0 & 0 & −I_{4} & 0 \\ 0 & 0 & 0 & −I_{4} \end{matrix} )

γ^{4+j} = E^{(4+j)} ⊗ I_{4} = −E^{j} ⊗ I_{4} = −γ^{j}, j = 1, 2, 3, 4.

So for this example, G_{8} = {γ^{i} = E^{i} ⊗ I_{4}: i = 1, 2, …, 8} is the cyclic subgroup of SP_{16} of order 8.

Lemma 7 Let q = (2n−1)^{2} be a prime power. Then there is a balanced weighing matrix BGW (q^{m} + q^{m−1} + ⋯ + q + 1, q^{m}, q^{m} − q^{m−1}) over the cyclic group G_{4n} for each positive integer m.

Proof.

Note that 4n is a divisor of q − 1 and apply [2], IV.4.22.

4 Symmetric designs with the Ionin–type parameters

We are now ready for the main result of the paper.

Theorem 8 Let H be a Bush-type Hadamard matrix of order 4n^{2} and q = (2n − 1)^{2} a prime power. Then there is a weighing matrix

W (4(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, 4q^{m}n^{2})

which contains a twin design with Ionin–type parameters,

ν_{m} = 4(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, κ_{m} = q^{m}(2n^{2} − n), λ_{m} = q^{m}(n^{2} − n),

for each positive integer m.

Proof.

Let m be a positive integer. Let W = [w_{ij}] be the balanced generalized weighing matrix BGW (ν, κ, λ) of the lemma 7, where ν = q^{m} +q^{m−1} +⋯+q +1, κ = q^{m}, λ = q^{m} −q^{m−1}. Consider the block matrix A = [Hw_{ij}] of order 4νn^{2}. Let AA^{t} = [B_{kl}].

For k ≠ l,

B_{kl} = ∑_{j=1}^{ν} Hw_{kj}(Hw_{jl})^{t}
= ∑_{j=1}^{ν} H(w_{kj}w_{lj}^{t})H^{t}
= H(∑_{j=1}^{ν} w_{kj}w_{lj}^{t}) H^{t}
= H( ∑_{γ∈G_{4n}} \frac{λ}{4n}γ) H^{t} = O.

For k = l,

B_{kl} = H ∑_{j=1}^{ν}(w_{kj}w_{lj}^{t}) H^{t}
= κHH^{t}
= 4n^{2}κI_{2n}.

So A is a weighing matrix W (4n^{2}ν, 4n^{2}κ). Now, let D = [Mw_{ij}], where M is the matrix of lemma 3, obtained from H by replacing all the diagonal blocks with the zero matrix.

This matrix contains two symmetric designs with the Ionin–type parameters. To see this, let

D = D^{+} − D^{−} = [Mw_{ij}].

Now note that M = H − I_{2n} ⊗ J_{2n} and it is easy to see that

D^{+} + D^{−} = [P |w_{ij}|].

Therefore we have,

2D^{+} = [P |w_{ij}| + Mw_{ij}].

First note that for every i, j, k, l, (P |w_{ij}|)(Mw_{kl})^{t} = (Mw_{kl})(P |w_{ij}|)^{t} = 0. Also note that a calculation on D = [Mw_{ij}] akin of that on A above would give,

[Mw_{ij}][Mw_{ij}]^{t} = κ(I_{ν} ⊗ MM^{t}).

So, we have,

4D^{+}D^{+t} = [P |w_{ij|}][P |w_{ij}|]^{t} + [Mw_{ij}][Mw_{ij}]^{t}
= [P |w_{ij}|][P |w_{ij}|]^{t} + κ(I_{ν} ⊗ MM^{t}).

For k ≠ l,

∑_{j=1}^{ν} P |w_{kj}|(P |w_{jl}|)^{t} = ∑_{j=1}^{ν} P (|w_{kj}||w_{lj}|^{t})P^{t}
= \frac{λ}{\frac{q−1}{2(n−1)}} P (J_{2n} ⊗ I_{2n})P^{t}
= \frac{2λ(n − 1)}{q − 1}((J_{2n} − I_{2n}) ⊗ J_{2n})(J_{2n} ⊗ I_{2n})((J_{2n} − I_{2n}) ⊗ J_{2n})
= \frac{2λ(n − 1)}{q − 1}2n(2n − 1)^{2}J_{2n} ⊗ J_{2n}
= \frac{2n − 1}{q − 1}2nq^{m−1}(q − 1)qJ_{2n} ⊗ J_{2n}
= 4n(n − 1)q^{m}J_{2n} ⊗ J_{2n}.

Therefore, all the (k, l), 1 ≤ k ≠ l ≤ ν blocks of the matrix D^{+}D^{+t} consist of the matrix n(n − 1)q^{m}J_{2n} ⊗ J_{2n}. For k = l, it follows from the alternate proof of lemma 3 that,

∑_{j=1}^{ν} P |w_{kj}|(P |w_{kj}|)^{t} + ∑_{j=1}^{ν} Mw_{kj}(Mw_{kj})^{t} = κ(P P^{t} + MM^{t})
= q^{m}(4n^{2}I_{4n^{2}} + 4n(n − 1)J_{2n} ⊗ J_{2n}).

From this we conclude that all the (k, k), 1 ≤ k ≤ ν blocks of the matrix D^{+}D^{+t} consist of the matrix q^{m}(n^{2}I_{4n^{2}} +n(n−1)J_{2n}⊗J_{2n}). Therefore, D^{+} is a symmetric (ν_{m}, κ_{m}, λ_{m})– design.

Noting that,

2D^{−} = [P |w_{ij}| − Mw_{ij}]

an almost identical argument shows that D^{−} is also a symmetric design with Ionin–type parameters.

Corollary 9 Let 4n be the order of an Hadamard matrix with q = (4n − 1)^{2} a prime power. Then there is a weighing matrix W (16(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, 16q^{m}n^{2}) which includes two symmetric designs with the Ionin–type parameters

ν = 16(q^{m} + q^{m−1} + ⋯ + q + 1)n^{2}, κ = q^{m}(8n^{2} − 2n), λ = q^{m}(4n^{2} − 2n)

for every positive integer m.

Proof.

This follows from theorems 1 and 8.

Remark 10

Corollary 9 provides a new class of symmetric designs. If a Bush–type Hadamard matrix of order 4n^{2} exists for odd values of n, then the construction methods in this paper simplifies Ionin’s method significantly. It is also interesting to note that there is a similarity between the way that we get twin designs from Bush–type Hadamard matrices and the way that the twin designs with the Ionin–type parameters are obtained from the weighing matrices.

Acknowledgment. The research is supported by an NSERC operating grant.

References

[1] K.A. Bush, Unbalanced Hadamard matrices and finite projective planes of even order, JCT, 11(1971), pp. 38–44.

[2] Warwick de Launey, Section on “Bhaskar Rao designs,” CRC Handbook of Combinatorial Designs, edited by Charles J. Colbourn and Jeffrey H. Dinitz, Kluwer Academic Press, 1996.

[3] Y.J. Ionin, New symmetric designs from regular Hadamard matrices, The Electronic Journal of Combinatorics, Vol. 5, No.1 (1998), R1.

[4] F. Kamali, H. Kharaghani and G. B. Khosrovshahi, Constructing some 2– (64,28,12) and 2–(144,66,30) symmetric designs, preprint.

[5] H. Kharaghani, New classes of weighing matrices, ARS comb., 19(1985), pp. 69–72.